OpenStudy (anonymous):

prove the trigonometric identities. sec θ / tan θ = csc θ Calculation Reason

Nnesha (nnesha):

rewrite sec and tan in terms of sin and cos theta

Nnesha (nnesha):

what's the definition of sec and tan ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i suck at trig

OpenStudy (anonymous):

plz help

Nnesha (nnesha):

cos ,sin ,tan and csc , sec , cot are reciprocal \[\cos \theta = \frac{1}{\sec \theta} ~~~~~\sin \theta=\frac{1}{\csc \theta} ~~~~\tan \theta =\frac{\sin \theta}{\cos \theta}\]

Nnesha (nnesha):

that's the definition you have to remember that btw \[\tan \theta = \frac{1}{\cot \theta} =\frac{\sin \theta}{\cos \theta}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so how do i solve the question

Nnesha (nnesha):

replace sec theta with 1/cos and tan with sin/cos

OpenStudy (anonymous):

i just need three Calculation Reason

Nnesha (nnesha):

\[\huge\rm \frac{ \sec \theta }{ \tan \theta }=\frac{\frac{1}{\cos \theta}}{\frac{\sin \theta}{\cos \theta}}\] now simplify

OpenStudy (anonymous):

im stuck

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha cos ,sin ,tan and csc , sec , cot are reciprocal \[\cos \theta = \frac{1}{\sec \theta} ~~~~~\sin \theta=\frac{1}{\csc \theta} ~~~~\tan \theta =\frac{\sin \theta}{\cos \theta}\] \(\color{blue}{\text{End of Quote}}\) these are the trig rules /reciprocal identities

Nnesha (nnesha):

here is an example how to simplify that `change division to multiplication` multiply top fraction by the reciprocal of the bottom fraction \[\large \rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } =\frac{ a }{ b } *\frac{ d }{ c }\]

OpenStudy (anonymous):

ok

Nnesha (nnesha):

\[\huge\rm \frac{ \sec \theta }{ \tan \theta }=\frac{\frac{1}{\cos \theta}}{\frac{\sin \theta}{\cos \theta}}\] try to simplify this by looking at that example let m know what u get

OpenStudy (anonymous):

theta=pi/2+2pin

Nnesha (nnesha):

hmm what ?? we don't need exact value we have to prove that L.H.S = R.H.S left side is equal to right

Nnesha (nnesha):

\[\huge\rm \frac{\frac{1}{\cos \theta}}{\frac{\sin \theta}{\cos \theta}}\] how would you change division to multiplication ?

OpenStudy (anonymous):

subtract from both sides

Nnesha (nnesha):

we are working on left side rewrite sec and tan in terms of cos and sin we got this \[\huge\rm \frac{\frac{1}{\cos \theta}}{\frac{\sin \theta}{\cos \theta}}\] now change division to multiplication we just working one one side

Nnesha (nnesha):

here is an example `change division to multiplication` multiply top fraction by the reciprocal of the bottom fraction \[\large \rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } =\frac{ a }{ b } *\frac{ d }{ c }\]

OpenStudy (anonymous):

0

Nnesha (nnesha):

how did you get 0 show ur steps

OpenStudy (anonymous):

multiplied it

Nnesha (nnesha):

what did you multiply ?

OpenStudy (anonymous):

a/b*d/c

Nnesha (nnesha):

right now how would you convert this division \[\huge\rm \frac{\frac{1}{\cos \theta}}{\frac{\sin \theta}{\cos \theta}}\] to multiplication

OpenStudy (anonymous):

i suck at trig baby steps plz

Nnesha (nnesha):

nnesha that's my username. anywayz forget about sin cos theta that's same as simple algebra let cos theta = x and sin theta = y so \[\large\rm \frac{ \frac{ 1 }{ x } }{ \frac{ y}{ x} }\]now can you simplify this ?

OpenStudy (anonymous):

is this relating to my question

Nnesha (nnesha):

yes

OpenStudy (anonymous):

i really dont know

Nnesha (nnesha):

well i don't know what you don't understand